OpenStudy (anonymous):

Can anyone help me with this? Find the solution to the following problem in correct scientific notation: (4.6 × 10^80) ÷ (2.3 × 10^40) A. 2.0 × 10^2 B. 2.0 × 10^40 C. 2.0 × 10^80 D. 2.0 × 10^120 E. 2.0 × 10^3,200

OpenStudy (mayankdevnani):

first, find the value of : \[\huge \frac{4.6 \times 10^{80}}{2.3 \times 10^{40}}\]

OpenStudy (mayankdevnani):

so what is the answer??? @RH

OpenStudy (mayankdevnani):

can i help you to find the answer of this fraction? @RH

OpenStudy (anonymous):

Yes if you don't mind :(

OpenStudy (mayankdevnani):

okkk.. so answer is : \[\huge 2.0 \times 10^{40}\]

OpenStudy (mayankdevnani):

here the proof of this answer: http://www.wolframalpha.com/input/?i=%7B+4.6+*+10+%5E+%7B+80+%7D+%7D+%2F+%7B+2.3+*+10+%5E+%7B+40+%7D+%7D

OpenStudy (anonymous):

Thank you!!! But I don't understand how to solve it... :(

OpenStudy (mayankdevnani):

okk... i will help you

OpenStudy (anonymous):

Thank you so much:(

OpenStudy (mayankdevnani):

suppose we have this value:  what is its decimal value

OpenStudy (mayankdevnani):

@RH can you write its value in decimal?

OpenStudy (anonymous):

0.002

OpenStudy (anonymous):

right?

OpenStudy (mayankdevnani):

right

OpenStudy (mayankdevnani):

and when i change this :  is there any affect on this value?